# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

### v2

@@ 1.3. $\mu$ transmission spaces
Here the boundary contributions are completely local. The formula was first proved for fractional laplacian in [Aba13] and later generalized for any classical pseudodifferential operator of order $2a$ in [Gru16a], [Gru20]. In articles [Gru16a], [Gru20] author shows that the boundary terms ${\partial_{\nu}\left(\frac{u}{d^{a-1}}\right),\partial_{\nu}\left(\frac{v}{d^{a-1}}\right)\in H^{s-a-\frac{1}{2}}(\partial\Omega)}$ (in case of (14) ${\frac{u}{d^{a}}\in H^{s-a-\frac{1}{2}}(\partial\Omega)}$ )))). Hence if $s>a+\frac{1}{2}$, then the $\mathrm{L}^{2}$ integral at the boundary makes sense. We would like to note that for any $a\in(\frac{1}{2},1)$ and for any $t>0$, the integration by parts formula holds for the functions $u_{f}$ and $u_{F}$, which are the solutions to (1) and (2), respectively, where $f\in\mathrm{L}^{2}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\partial\Omega)$. Thanks to theorem 7 and theorem 8, the formula clearly holds for $f\in C_{c}^{\infty}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\Omega)$. Let $f\in\mathrm{L}^{2}(\Omega)$, then there exists a sequence $f_{m}\in C_{c}^{\infty}(\Omega)$ such that $f_{m}\to f$ in $\mathrm{L}^{2}(\Omega)$. For all $m\in\mathbb{N}$, we have
(15) $\int_{\Omega}u_{f_{m}}(t,\cdot)(-\Delta)^{a}u_{F}(t,\cdot)-$ $u_{F}(t,\cdot)(-\Delta)^{a}u_{f_{m}}(t,\cdot)$
$=$ $-\Gamma(a)\Gamma(a+1)\int_{\Omega}\frac{u_{F}}{d^{a-1}}(t,\cdot)\frac{u_{f_{m}}}{d^{a}}(t,\cdot).$
Thanks to [CGCV22], we can represent the solution in the following way:
$u_{f_{m}}(t,\cdot)=\sum\limits_{n\in\mathbb{N}}e^{-t\lambda_{n}}\phi_{n}(x)\langle\phi_{n},f_{m}\rangle,$
where $\phi_{n}$’s are the eigenfunctions corresponding to eigenvalue $\lambda_{n}$, satisfying (6). For a given $t>0$, we have $u_{f_{m}}(t,\cdot)\to u_{f}(t,\cdot)$, $(-\Delta)^{a}u_{f_{m}}(t,\cdot)\to(-\Delta)^{a}u_{f}(t,\cdot)$ and ${\frac{u_{f_{m}}}{d^{a}}(t,\cdot)\to\frac{u_{f}}{d^{a}}(t,\cdot)}$ in $\mathrm{L}^{2}(\Omega)$. Hence, taking $m\to+\infty$ in (15), we obtain
(16) $\int_{\Omega}u_{f}(t,\cdot)(-\Delta)^{a}u_{F}(t,\cdot)-$ $u_{F}(t,\cdot)(-\Delta)^{a}u_{f}(t,\cdot)$
$=$ $-\Gamma(a)\Gamma(a+1)\int_{\Omega}\frac{u_{F}}{d^{a-1}}(t,\cdot)\frac{u_{f}}{d^{a}}(t,\cdot),$
for any $t\in(0,T]$.
We also mention a particular integration by parts formula which we will use throughout the article. Let $a\in(0,1)$ and $w_{1},w_{2}\in H^{a}(\mathbb{R}^{N})$ with $w_{1}\equiv w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$, then
@@
This can be established easily by considering the integral in Fourier space.
@@ 2. Density Results:
Let $u_{f^{*}}\in\mathrm{L^{2}((0,\tilde{T});H^{a}(\Omega)})$ be the solution of the adjoint system (13) with $f^{*}$ as initial data. We use the integration by parts formula (17)
@@
$\int_{\Omega}u_{f^{*}}(\tilde{T},\cdot)u_{f}(\tilde{T},\cdot)=\int_{\Omega}u_{f^{*}}(0,\cdot)u_{f}(0,\cdot).$
Thanks to (18) and since $f\in C_{c}^{\infty}(\Omega)$ is arbitrary, we deduce that
@@
Furthermore, the solution can be represented in the following form [see [CGCV22]]:
@@
where for all $n\in\mathbb{N}$, $\phi_{n}$’s are the eigenfunctions corresponding to (6). We have that $\lambda_{1}<\lambda_{2}\leq\dots\lambda_{n}\leq\dots+\infty$ and $\{\phi_{n}\}$ is an orthonormal basis of $\mathrm{L}^{2}(\Omega)$. Hence $u_{f^{*}}(0,x)\equiv 0$ implies
@@
This further implies $f^{*}\equiv 0$, contradicts $\|f^{*}\|_{\mathrm{L}^{2}(\Omega)}=1$. Hence, for a given $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$.
Second proof: We now provide a quantitative proof based on variational approach.
@@
    Continuity: Let $f_{n}^{*}\to f^{*}$ in $\mathrm{L}^{2}(\Omega)$. The solution of (13), $u_{f^{*}_{m}}$ corresponding to the initial data $f_{m}^{*}$ can be expressed as
@@
    where $\phi_{n}$ are the eigenfunctions of (6). Hence
@@
    Coercivity: Let $f_{m}^{*}\in\mathrm{L}^{2}(\Omega)$, for all $m\in\mathbb{N}$ and ${\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ as $m\to+\infty$. Consider the functions ${g_{m}^{*}=\frac{f_{m}^{*}}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. Then ${\|g_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}=1}$ for all $m\in\mathbb{N}$. Let $u_{g^{*}_{m}}$ be the solution of (13) with respect to the initial data $g_{m}^{*}$. Then ${u_{g^{*}_{m}}=\frac{u_{f^{*}_{m}}}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. We divide the functional $\mathcal{J}_{\epsilon,a}$ by $\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}$. It yields
@@
    where $\phi_{n}$’s are the eigenfunctions corresponding to (6). Let $u_{g^{*}}$ be the solution of (1) with initial data $g^{*}$.
@@
    $+$ $4C_{1}(a)\sum\limits_{n=n_{K}+1}^{\infty}\frac{1}{\lambda_{k}^{\frac{n}{s}}}$
@@
Continuity, strict convexity and coercivity helps us conclude that $\mathcal{J}_{\epsilon,a}$ attains it’s minimum for some $f_{min}^{*}\in\mathrm{L}^{2}(\Omega)$. We fix a function $g^{*}\in\mathrm{L}^{2}(\Omega)$ and let $u_{g^{*}}$ be the solution of (13) corresponding to the initial data $g^{*}$. We have
@@
Hence for a given $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$. ∎
Next we are interested in singular boundary data system (2). We like to analyze the density in $\mathrm{L}^{2}(\Omega)$ space on a particular time slice by varying singular boundary data. We start with an observability estimate.
@@
Let $f^{*}\in L^{2}(\Omega)$ and let $0\leq t_{1}<t_{2}\leq T$. Let $u_{f^{*}}$ be the solution of (13) with initial data $f^{*}$. Then the following observability estimate holds
$\int_{\Omega}u_{f^{*}}^{2}(t_{1},\cdot)\,\rm{d}x-$ $\int_{\Omega}u_{f^{*}}^{2}(t_{2},\cdot)\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x\,\rm{d}t$
$=$ $-\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}t.$
Furthermore there exists a positive constant $\mathcal{A}$, depends only on the domain, the dimension and the exponent ‘$a$’, such that:
$\mathcal{A}\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}s\geq\|u_{f^{*}}\|_{\mathrm{L}^{2}((t_{1},t_{2})\times\Omega)}.$
@@
Thanks to (9), we have $\nabla u_{f^{*}},(-\Delta)^{a}u_{f^{*}}\in\mathrm{L}^{2}(\Omega)$ for all $t\in(0,T)$ and thanks to (7),(8),(10),(11) we can apply fractional Pohozaev identity (12) for the function $u_{f^{*}}$. It yields
$\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $u_{f^{*}}(-\Delta)^{a}u_{f^{*}}\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
Equation (13) yields
(24) $\int_{\Omega}(x\cdot\nabla u_{f^{*}})$ $(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x$
-- $\frac{2a-n}{2}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
We intend to calculate the term ${\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x}$. Gauss divergence theorem yields:
$\int_{\Omega}(x\cdot\nabla u_{f^{*}})$ $(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=-\int_{\Omega}\nabla\cdot\left(x(-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x$
$=$ $-\int_{\Omega}n(-\Delta)^{a}u_{f^{*}}u_{f^{*}}\,\rm{d}x-\int_{\Omega}x\cdot\nabla\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x.$
Using equation (13), we obtain
(25) $\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=$ $-n\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x+n\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x$
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x.$
We compute the last term of the right hand side above:
$-\int_{\mathbb{R}^{N}}x\cdot\nabla$ $\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x=-\int_{\mathbb{R}^{N}}\nabla_{\xi}\cdot\left(\xi|\xi|^{2a}\hat{u_{f^{*}}}\right)\hat{u_{f^{*}}}\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}\left(n|\xi|^{2a}\hat{u_{f^{*}}}+2a|\xi|^{2a}\hat{u_{f^{*}}}+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\hat{u_{f^{*}}}\right)\right)\hat{u_{f^{*}}}\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}n|\xi|^{2a}\hat{u_{f^{*}}}^{2}+2a|\xi|^{2a}\hat{u_{f^{*}}}^{2}+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\frac{\hat{u_{f^{*}}}^{2}}{2}\right)\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}(n+2a)|\xi|^{2a}\hat{u_{f^{*}}}^{2}\,\rm{d}\xi+\int_{\mathbb{R}^{N}}\nabla_{\xi}.\left(\xi|\xi|^{2a}\right)\frac{\hat{u_{f^{*}}}^{2}}{2}\,\rm{d}\xi$
$=$ $-\frac{n+2a}{2}\int_{\mathbb{R}^{N}}|\xi|^{2a}\hat{u_{f^{*}}}^{2}\,\rm{d}\xi=-\frac{n+2a}{2}\int_{\Omega}(-\Delta)^{a}u_{f^{*}}u_{f^{*}}\,\rm{d}x$
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}u_{f^{*}}u_{f^{*}}-qu_{f^{*}}^{2}\right)\,\rm{d}x$
$=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)-qu_{f^{*}}^{2}\right)\,\rm{d}x.$
Substituting this in (25), we obtain
$\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=-\frac{3n+2a}{2}\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x+\frac{3n+2a}{2}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x.$
Substituting this in (24) yields
$(n+2a)\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x-(n+2a)\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma=0.$
Integrating with respect to time yields:
(26) $\int_{\Omega}u_{f^{*}}^{2}(t_{1},\cdot)\,\rm{d}x-$ $\int_{\Omega}u_{f^{*}}^{2}(t_{2},\cdot)\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x\,\rm{d}t$
$=$ $-\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}t.$
We recall equation (13):
(27) $-\partial_{t}u_{f^{*}}+(-\Delta)^{a}u_{f^{*}}+qu_{f^{*}}=0$
with $u_{f^{*}}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$. We multiply the equation (27) by $u_{f^{*}}$ and integrate with respect to space time.
$\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{2},\cdot)\,\rm{d}x+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}u_{f*}$ $(-\Delta)^{a}u_{f^{*}}\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}qu^{2}_{f*}\,\rm{d}x\,\rm{d}t=0.$
This implies
$\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{2},\cdot)\,\rm{d}x+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}$ $\left|(-\Delta)^{\frac{a}{2}}u_{f^{*}}\right|^{2}\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}qu^{2}_{f*}\,\rm{d}x\,\rm{d}t=0,$
i.e.
$\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{2},\cdot)\,\rm{d}x+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}qu^{2}_{f*}\,\rm{d}x\,\rm{d}t$
$=-\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}u_{f^{*}}\right|^{2}\,\rm{d}x\,\rm{d}t.$
By Hardy-Littlewood-Sobolev estimate:
$\|u_{f^{*}}\|_{\mathrm{L}^{2}(\Omega)}\leq C_{Vol}\|u_{f^{*}}\|_{\mathrm{L}^{\frac{2n}{n-a}}(\Omega)}\leq C_{n,a,\Omega}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}u_{f^{*}}\right|^{2}\,\rm{d}x\,\rm{d}t,$
where $C_{Vol}$ and $C_{n,a,\Omega}$ are two fixed positive constants depends on the domain and the dimension only. Let ${\mathcal{A}=\frac{1}{2C_{n,a,\Omega}}}$. Then from (26), we conclude
$\mathcal{A}\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}s\geq\|u_{f^{*}}\|_{\mathrm{L}^{2}((t_{1},t_{2})\times\Omega)}.$
We now move onto the proof of our next density result.
@@
Again we present two proofs here. One using integration by parts (14) and another by a variational approach.
First proof: Let the set ${\{u_{F}(\tilde{T},\cdot),\,F\in C_{c}^{\infty}((0,T)\times\Omega)\}}$ is not dense in $\mathrm{L}^{2}(\Omega)$. Then by Hahn-Banach theorem there exists $g\in\mathrm{L}^{2}(\Omega)$ such that $\|g\|_{L^{2}(\Omega)}=1$ and
(28) $\langle u_{F}(\tilde{T},\cdot),g\rangle=0,\qquad\forall F\in C_{c}^{\infty}((0,T)\times\Omega).$
Let $u_{g}\in\mathrm{L^{2}((0,\tilde{T});H^{a}(\Omega))}$ be the solution of the adjoint system (13) with $g$ as initial data at the time $T=\tilde{T}$. Theorem 8 ensures ${u_{F}\in\mathrm{L}^{2}\left((0,\tilde{T});H^{(a-1)(2a)}(\overline{\Omega})\right)}$ respectively. We apply integration by parts formula (16).
@@
$\int_{\Omega}\partial_{t}u_{g}u_{F}+u_{g}\partial_{t}u_{F}=\int_{\Omega}\partial_{t}\left(u_{g}u_{F}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{a}}.$
Integrating with respect to $t$ yields:
@@
Thanks to (28) and since $F\in C_{c}^{\infty}((0,T)\times\Omega)$ is arbitrary, we deduce that
@@
Observability estimate (9) yields
$\|u_{g}\|_{\mathrm{L}^{2}((0,\tilde{T})\times\Omega)}=0\implies u_{g}\equiv 0\implies g\equiv 0.$
Next we present a proof via variational approach. It’s more quantitative in nature.

### v3

@@ 1.4. Boundary values of $\mu$ transmission spaces, Dirichlet trace
Here the boundary contributions are completely local. The formula was first proved for fractional Laplacian in [Aba13] and later generalized for any classical pseudodifferential operator of order $2a$ in [Gru16a], [Gru20]. In articles [Gru16a], [Gru20], author shows that the boundary terms ${\partial_{\nu}\left(\frac{u}{d^{a-1}}\right),\partial_{\nu}\left(\frac{v}{d^{a-1}}\right)\in H^{s-a-\frac{1}{2}}(\partial\Omega)}$ (in case of (16) ${\frac{u}{d^{a}}\in H^{s-a-\frac{1}{2}}(\partial\Omega)}$ )). Hence if $s>a+\frac{1}{2}$, then the $\mathrm{L}^{2}$ integral at the boundary makes sense. We would like to note that for any $a\in(\frac{1}{2},1)$ and for any $t>0$, the integration by parts formula holds for the functions $u_{f}$ and $u_{F}$, which are the solutions to (1) and (2), respectively, where $f\in\mathrm{L}^{2}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\partial\Omega)$. thanks to theorem 7 and theorem 8.
We also mention a particular integration by parts formula which we will use throughout the article. Let $a\in(0,1)$ and $w_{1},w_{2}\in H^{a(2a)}(\overline{\Omega})\cap H^{a}(\mathbb{R}^{N})$ with $w_{1}\equiv w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$, then
@@
This can be established easily using linear translation.
@@ 2. Density Results:
This section we start with the density of the solution space corresponding to (1).
@@
Let $u_{f^{*}}$ be the solution of the adjoint system (14) with $f^{*}$ as initial data. Thanks to Theorem 7, we have that
$u_{f^{*}}\in C\left((0,\tilde{T});H^{a(2a)}(\overline{\Omega})\right)\cap C^{1}((0,\tilde{T});\mathrm{L}^{2}(\Omega))\cap C([0,\tilde{T}];\mathrm{L}^{2}(\Omega)).$
We use the integration by parts formula (17)
@@
$\int_{\Omega}u_{f^{*}}(\tilde{T}-\epsilon,\cdot)u_{f}(\tilde{T}-\epsilon,\cdot)=\int_{\Omega}u_{f^{*}}(0,\cdot)u_{f}(0,\cdot),\quad\forall\,\epsilon>0.$
Thanks to continuity with respect to time variable (Theorem 7) and (18), we deduce that
@@
Since $f\in C_{c}^{\infty}(\Omega)$ is arbitrary. Furthermore, the solution can be represented in the following form [see [CGCV22]]:
@@
where for all $n\in\mathbb{N}$, $\phi_{n}$’s are the normalized eigenfunctions corresponding to (6). We have that $\lambda_{1}<\lambda_{2}\leq\dots\lambda_{n}\leq\dots+\infty$ and $\{\phi_{n}\}$ is an orthonormal basis of $\mathrm{L}^{2}(\Omega)$. Hence $u_{f^{*}}(0,x)\equiv 0$ implies
@@
This further implies $f^{*}\equiv 0$, contradicts $\|f^{*}\|_{\mathrm{L}^{2}(\Omega)}=1$. Hence, for a given $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $\mathrm{L}^{2}(\Omega)$.
Second proof: We now provide a quantitative proof based on variational approach. We prove that the solution set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $\mathrm{L}^{2}(\Omega)$.
@@
    Continuity: Let $f_{n}^{*}\to f^{*}$ in $\mathrm{L}^{2}(\Omega)$. The solution to (14), $u_{f^{*}_{m}}$ corresponding to the initial data $f_{m}^{*}$ can be expressed as
@@
    where $\phi_{n}$ are the normalized eigenfunctions of (6). Hence
@@
    Coercivity: Let $f_{m}^{*}\in\mathrm{L}^{2}(\Omega)$, for all $m\in\mathbb{N}$ and ${\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ as $m\to+\infty$. Consider the functions ${g_{m}^{*}=\frac{f_{m}^{*}}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. Then ${\|g_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}=1}$ for all $m\in\mathbb{N}$. Let $u_{g^{*}_{m}}$ be the solution to (14) with respect to the initial data $g_{m}^{*}$. Then ${u_{g^{*}_{m}}=\frac{u_{f^{*}_{m}}}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. We divide the functional $\mathcal{J}_{\epsilon,a}$ by $\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}$. It yields
@@
    where $\phi_{n}$’s are the eigenfunctions corresponding to (6). Let $u_{g^{*}}$ be the solution to (1) with initial data $g^{*}$.
@@
    $+$ $4C_{1}(a)\sum\limits_{n=n_{K}+1}^{\infty}\frac{1}{|\lambda_{k}|^{\frac{N}{s}}}$
@@
Continuity, strict convexity and coercivity helps us conclude that $\mathcal{J}_{\epsilon,a}$ attains it’s minimum for some $f_{min}^{*}\in\mathrm{L}^{2}(\Omega)$. We fix a function $g^{*}\in\mathrm{L}^{2}(\Omega)$ and let $u_{g^{*}}$ be the solution to (14) corresponding to the initial data $g^{*}$. We have
@@
Hence for a given $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $\mathrm{L}^{2}(\Omega)$. ∎
Next we are interested in singular boundary data system (2). We like to analyze the density in $\mathrm{L}^{2}(\Omega)$ space on a particular time slice by varying singular boundary data. We describe two different ways to obtain the density of the solution space. One through observability estimate (see [BWZ25]) and the other through boundary unique continuation type result. However, the observability estimate only holds for small non-negative potential ‘$q$’ with small growth. We start with an observability estimate.
@@
Let $a\in\left(\frac{1}{2},1\right)$ and let $f^{*}\in\mathrm{L}^{2}(\Omega)$ and let $0\leq\epsilon<T$. Let $u_{f^{*}}$ be the solution to (14) with initial data $f^{*}$. Let ${\theta<\frac{1}{2}\left(1+C_{HS}\left(\frac{N}{2}+R\right)\right)^{-1}}$, where $C_{HS}$ is the Hardy-Sobolev constant on the domain $\Omega$ and $R:=\max\{\|x\|,x\in\Omega\}$. Let the potential $q$ is non-negative and $|q(x)|,|\nabla q(x)|\leq\theta$ for all $x\in\Omega$. Then the following observability estimate holds
$\mathcal{A}\int_{0}^{T-\epsilon}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}s\geq\|u_{f^{*}}(0)\|_{\mathrm{L}^{2}(\Omega)},$
where, the positive constant $\mathcal{A}$, depends only on the domain, the dimension, the time factor $T-\epsilon$ and the exponent ‘$a$’.
@@
The proof is exactly similar to the proof of observability estimate in [BWZ25]. The only difference is we have a small potential here, which was not present in [BWZ25]. The idea is to obtain a control function that renders the solution null controllable within a specific finite-dimensional subspace,
(24) $\mathcal{H}_{J}:=span\{\phi_{1},\cdots,\phi_{J}\},$
and then use this control as a test function in the adjoint system with the given initial data (14). Here $\{\phi_{i}\}\Big|_{i\in\mathbb{N}}$ is the eigenfunctions of the operator $(-\Delta)^{a}+q$, which generates an orthonormal basis of $\mathrm{L}^{2}(\Omega)$. In [BWZ25], the authors achieved this by making a detour through the wave equation, establishing an observability estimate for its energy, specifically for solutions lying in the finite-dimensional subspace $\mathcal{H}_{J}$. Although in [BWZ25] authors don’t have a potential ‘$q$’. We claim that even if we have a small potential we can obtain similar result. So, in order to prove Proposition 10, we make a little detour to the wave equation with a small potential and will obtain a observability estimate corresponding to it’s energy, when the solutions are coming from the space $\mathcal{H}_{J}$. We move the proof in Appendix-5. ∎
The above Proposition enables us to prove the density of the solution space corresponding to the singular boundary data (2).
@@
Again we present two proofs here. One using integration by parts (16) and another by a variational approach. Due to technical reason, for the variational approach, we will assume the potential is small, non-negative and it’s space growth is also small.
First proof: Let the set ${\{u_{F}(\tilde{T},\cdot),\,F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ is not dense in $\mathrm{L}^{2}(\Omega)$. Then, by Hahn-Banach theorem, there exists $g\neq 0$ in the normal of the closure of the solution set ${\{u_{F}(\tilde{T},\cdot),\,F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ such that $\|g\|_{\mathrm{L}^{2}(\Omega)}=1$ and
(25) $\langle u_{F}(\tilde{T},\cdot),g\rangle=0,\qquad\forall F\in C_{c}^{\infty}((0,T)\times\partial\Omega).$
Let $u_{g}$ be the solution to the adjoint system (14) with $g$ as the initial data at the time $T=\tilde{T}$. Thanks to Theorem 7, we have that
$u_{g}\in C\left((0,\tilde{T});H^{a(2a)}(\overline{\Omega})\right)\cap C^{1}((0,\tilde{T});\mathrm{L}^{2}(\Omega))\cap C([0,\tilde{T}];\mathrm{L}^{2}(\Omega)).$
We apply integration by parts formula (16):
@@
$\int_{\Omega}u_{F}\partial_{t}u_{g}+u_{g}\partial_{t}u_{F}=\int_{\Omega}\partial_{t}\left(u_{g}u_{F}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{a}}.$
Integrating with respect to $t$ and continuity in time variable yields:
@@
Thanks to (25) and since $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ is arbitrary, we deduce that
@@
Proposition 10 yields
$g\equiv 0\ \ \ \mbox{a contradiction}.$
Next we present a proof via variational approach. It is more quantitative in nature. We prove that the solution set ${\{u_{F}(\tilde{T},\cdot),\,F\in\mathrm{L}^{2}((0,T)\times\partial\Omega),F(0,\cdot)=0\}}$ is dense in $\mathrm{L}^{2}(\Omega)$. The theorem then follows because of the space $C_{c}^{\infty}((0,T)\times\Omega)$ is dense in $\mathrm{L}^{2}((0,T)\times\Omega)$ and [BWZ25][Proposition A.13]. In order to apply observability estimate we assume the potential is non-negative and small as described in Proposition 10.